OpenStudy (anonymous):

Which is the most reasonable estimate for 25% of 817?

OpenStudy (lgbasallote):

\[\frac{25}{100} \times 817 \implies \frac 14 \times 817 \implies 817 \div 4 \approx 204.25\] doess that help?

OpenStudy (anonymous):

yes it does thank you

OpenStudy (anonymous):

i think its more of a mental math type of problem round 817 down to 800 then find 25% of 800